OpenStudy (jpsmarinho):

How to integrate this: ln(2x+1) dx. I tried use integration by parts. But unsuccessful.

OpenStudy (amistre64):

ln(u) integrates to u ln(u) - u by tables :)

OpenStudy (amistre64):

integration by parts does work it out; you just have to use 1 dx your "dv"

OpenStudy (amistre64):

\[\int ln(u)du=u~lnu-\int \frac{1}{u}u~du\]

OpenStudy (jpsmarinho):

But, I go to this integral: X/(2x+1)dx and I not know how to integrate this ^^

OpenStudy (anonymous):

\[\int \frac{2x}{2x+1} \text{d}x\]?

OpenStudy (jpsmarinho):

Yes

OpenStudy (anonymous):

\[ \frac{2x}{2x+1}=\frac{2x+1-1}{2x+1}=1-\frac{1}{2x+1} \]does that help?

OpenStudy (jpsmarinho):

Thanks! I had forgot to use this

OpenStudy (anonymous):

np :)